OpenStudy (anonymous):

What Is the smallest multiple of 9, all of whose digits are even?

OpenStudy (tkhunny):

90 - How were you trying to solve this one?

OpenStudy (turingtest):

9 is not an even digit

OpenStudy (anonymous):

I don't understand how it could have an even multiple?

OpenStudy (tkhunny):

Yikes. I got all excited with the zero. Very bad. The questino stands. How ere you trying to solve this one? Can it be solved?

OpenStudy (turingtest):

@buttersnap 9*2=18 but 1 is odd 9*3 is 27 but 7 is odd find a multiple that is not odd

OpenStudy (anonymous):

Generated them Seems to be 288

OpenStudy (turingtest):

that seems to be the problem, I have no ideas how to approach it yet

OpenStudy (turingtest):

great, but I still want a method :P

OpenStudy (turingtest):

You could write a 3 line python program to solve it, but I don't think that's hte point here

OpenStudy (anonymous):

Exactly what I did lol. Wanted to make sure it happens If I remember correctly test for divisibility by 9 has to do with sum of digits.

OpenStudy (anonymous):

Thanks for all the help guys!

OpenStudy (tkhunny):

If there is a 3-digit solution, we have n*9 = 100*2j + 10*2k + 2l, where j, k, and l are digits less than 5. n must be a multiply of 2. It's not much, but it's something. 2j + 2k + 2l must be divisible by 3, so j+k+l must be 3, 6, 9, or 12. 2j + 2k + 2l must be divisible by 9, so j+k+l must be 9. There are hints, but it's not clean. I'll quit typing, now, in case I get all exicted about zero, again.

OpenStudy (anonymous):

I ignored 0. Might have been a trick question

Directrix (directrix):

288 may work.